Model A1311 / Late 2009 / 3.06 or 3.33 GHz Core 2 Duo processor

199 Questions View all

cpu fan speed 3975, optical 3800, HDD 1200

Hello, I've swapped my optical for an SSD, and replaced the HDD with a 2TB. I've found a second optical temp sensor and it seems to work fine on the replacement HDD. But the optical temp sensor reads zero, with fan speed 3800, and the cpu temp when reading 21, has fan speed of 3975.

Can anyone help with this fan speed issue?

Thanks

Answered! View the answer I have this problem too

Is this a good question?

Score 0
Add a comment

2 Answers

Chosen Solution

I found the problem. I haven't yet connected the LCD temp sensor. I have been using the external monitor connection to test both drives without hooking up the LCD. So, I think that's it. Thanks.

BTW the program SSDFanControll 2.1 was very helpful for controling the optical fan. The cpu fan speed was confusing me.

Was this answer helpful?

Score 0
Add a comment
Most Helpful Answer

The late 2009 Machine used Apple proprietary hard drives with special heat sensors. Most of the "off the shelf" hard drives will cause fan problems.

Please give us the last three figures of your serial number so we know your exact machine. Also exactly which hard drive that you had.

Was this answer helpful?

Score 1

Comments:

Thanks for the response.

The HDD fan seems correct at 1200. I used

a second optical sensor for the optical drive

and if you remove this sensor the HDD fan

will increase to 4250, but plugging it back

in brings the HDD fan back to 1200. I've

tested both optical sensors, with the same

result.

iMac serial -- V10

Seagate 1TB

(let me know what else you need here)

iMac 27 late 2009,

SMC ver 1.53f13 (smc 1.0)

by

iMac serial number last 3 -- V10 original HDD seagate 1TB

by

late 2009 27" iMac fans and temp sensors

3 fans, CPU, HDD, Optical

3 sensors LCD, HDD, Optical

Removing the LCD TS drives the CPU and Optical fans high

Removing the Opt TS drives only the Optical fan high,

same for HDD.

If either the LCD or Optical fan get driven high, the sensor

must be installed and the SMC reset to get the fans back

to normal.

Seems like all the sensors are the same.

by

Add a comment

Add your answer

stone will be eternally grateful.
View Statistics:

Past 24 Hours: 0

Past 7 Days: 0

Past 30 Days: 1

All Time: 452